# On the effects of similarity metrics in decentralized deep learning under distributional shift

### v1

@@ 1 Introduction
Among these approaches, federated learning (FL) has emerged as a leading framework for distributed deep learning, allowing the training of a global model across multiple clients that maintain private datasets (McMahan et al., 2017). Various studies have aimed to address client drift—a phenomenon occurring when clients pursue divergent optimization objectives due to distributional shifts between datasets, resulting in diminished performance upon averaging their model parameters. To counter this, several strategies introduce a regularization objective into the global training goal to minimize deviations between the global model and individual local models (Li et al., 2020; Karimireddy et al., 2020; Li et al., 2021). This approach is effective when clients share similar datasets, sampled from the same underlying data generating process. In these cases, a common global model is optimal. Conversely, in scenarios where clients’ data originate from distinct generating processes, these methods can be suboptimal and, in some cases, detrimental. Ghosh et al. (2020) address this issue within centralized federated learning by training multiple global models—specifically, $K$ models corresponding to the number of data clusters, with clustering determined by empirical loss to approximate data similarity. Similarly, Sattler et al. (2020) employ cosine similarity on local gradients for clustering clients in FL. Although effective, this leads to a high computational cost as the similarity is computed at stationary solutions of the FL objective.
@@ 1.1 Technical challenges and related work
One notable approach within decentralized learning is gossip learning, a peer-to-peer communication protocol, which has been extensively studied in various machine learning contexts (Kempe et al., 2003; Boyd et al., 2006; Ormándi et al., 2013; Hegedűs et al., 2019). Initial decentralized work using gossip-based optimization for non-convex deep learning tasks, specifically with convolutional neural networks (CNNs), was explored by Blot et al. (2016). However, gossip learning faces challenges in non independent and identically distributed (non-iid) settings where there is a distributional shift across client datasets. This is primarily due to the inadequacy of gossip learning in selectively identifying beneficial peers during the training process. To address these challenges, recent work has addressed decentralized learning within non-iid contexts. Onoszko et al. (2021); Li et al. (2022) introduced client selection algorithms based on empirical loss, employing distinct discovery and selections phases. These studies address challenges in decentralized learning within non-iid settings but require extensive hyperparameter tuning, particularly regarding search length, and rely on hard cluster assignments for clients. Such hard clustering can be detrimental if the discovery phase is noisy. To mitigate these issues, Listo Zec et al. (2022) proposed a sampling strategy that converts the similarity scores into sampling probabilities, which are used each communication round to sample collaborators. This approach identifies similar clients and facilitates model merging in proportion to their similarity. By adopting a soft clustering technique, it enhances the efficiency of peer collaboration, circumventing the need for the computationally intensive and difficult-to-tune discovery phase required by hard clustering methods. Additionally, Listo Zec et al. (2024) introduced a framework utilizing multi-armed bandits for client selection, where the selection process is modeled on empirical rewards derived from validation accuracy.
@@ 2 Problem formulation
Given that the true distribution $P(x,y)$ is unobserved, direct optimization of equation 1 is not possible. Instead, each client performs empirical risk minimization (ERM) locally by optimizing
$w^{\star}_{i}=\arg\min_{w_{i}\in\mathbb{R}^{d}}\mathbb{E}_{(x,y)\sim\mathcal{D}_{i}}\left[\ell(f_{i}(w_{i};x),y)\right].$ (2)
@@ 3 Experimental setup
In this section, we describe our experimental setup for studying the effect of common similarity metrics on client identification within decentralized learning. For reproducibility, our code is included in the supplementary materials and will be released on GitHub. Empirical loss and cosine similarity on gradients have been previously employed (Onoszko et al., 2021; Li et al., 2022; Listo Zec et al., 2022). Additionally, we introduce the Euclidean distance ($L^{2}$) between model parameters and cosine similarity between parameters (as opposed to gradients) as two additional baselines, which have not been tested before in DL for client identification. We define the similarity metrics between two clients $i$ and $j$ as follows:
@@ 3.2 Model selection and hyperparameter tuning
In both federated and decentralized learning environments, the process of model selection and the setting of hyperparameters are crucial yet often under-discussed elements. A common challenge in the literature is the lack of detailed reporting on hyperparameter choices in decentralized settings, which complicates the task of conducting fair and consistent methodological comparisons. Variations in results that stem from differing hyperparameter tuning practices can be mistakenly attributed to the intrinsic qualities of the algorithms themselves, thereby obscuring accurate evaluations. In practice, while some studies utilize a global validation set, others may employ a local validation set specific to each client, and still others might not clearly disclose their hyperparameter selection strategy.
In our investigation, we strive for a rigorous comparison across all models and similarity metrics by utilizing a local validation set for each client. We tune all hyperparameters using these validation sets, and employ early stopping on each client locally. During communication, some clients may reach early stopping before others. When this occurs, these clients continue to broadcast their models to other clients but cease updating their own models through further training or merging. This approach ensures that our evaluation is both relevant and tailored to the specific conditions of each client’s data. Additionally, we introduce a comparison against an Oracle baseline, which has access to clients with the same data distribution and only communicates with those, providing a benchmark under ideal conditions, and a Local baseline which only performs local training of the available data on a single client.
@@ 3.3 Distributional shifts and datasets
Label shift. Using the CIFAR-10 dataset (Krizhevsky et al., 2009), we create two more realistic and complex scenarios. The first scenario divides the dataset into two clusters based on content categories: an animal cluster, comprising four labels, and a vehicle cluster, with six labels. These clusters consist of 40 and 60 clients, respectively. In the second scenario, we create five equally large clusters of 20 clients where each cluster is defined by two unique, randomly sampled labels. In this way, each cluster has a unique data generating process where $P(y)$ varies while $P(x|y)$ remains constant. We also study large scale pre-trained models under label shift. For this, we utilize the CIFAR-100 dataset (Krizhevsky et al., 2009), which we split into three distinct clusters of sizes 26, 13 and 13 based on the superclasses.
@@ 4 Results
$L^{2}$ FedAvg 21.10 59.44 84.04 90.49 84.17 55.53
$L^{2}$ FedSim 10.85 57.63 84.06 90.80 84.16 -
Inv loss FedAvg 31.69 61.62 84.94 89.83 84.83 54.23
Inv loss FedSim 14.82 61.08 85.8 90.67 85.08 -
Cos grad FedAvg 10.32 58.81 86.22 91.14 84.87 53.87
Cos grad FedSim 10.30 58.13 86.0 90.72 83.32 -
Cos weight FedAvg 10.34 59.14 85.85 91.03 85.04 54.54
Cos weight FedSim 10.30 59.66 85.66 90.62 84.82 -
Random 1494.84 59.71 82.73 90.53 83.70 54.88
Local 30.26 37.77 77.08 82.22 76.26 29.49
@@ 6 Future work
Additionally, investigating the intersection of similarity metrics and privacy is a compelling direction. Key questions include: How robust are similarity metrics when subjected to privacy-preserving techniques such as differential privacy? Exploring these areas will deepen our understanding of decentralized learning and enhance the applicability and robustness of similarity metrics in diverse and privacy-sensitive settings.

### v2

@@ 1 Introduction
Among these approaches, federated learning (FL) has emerged as a leading framework for distributed deep learning, allowing the training of a global model across multiple clients that maintain private datasets (McMahan et al., 2017). Various studies have aimed to address client drift—a phenomenon occurring when clients pursue divergent optimization objectives due to distributional shifts between datasets, resulting in diminished performance upon averaging their model parameters. To counter this, several strategies introduce a regularization objective into the global training goal to minimize deviations between the global model and individual local models (Li et al., 2020; Karimireddy et al., 2020; Li et al., 2021). This approach is effective when clients share similar datasets, sampled from the same underlying data generating process. In these cases, a common global model is optimal. Conversely, in scenarios where clients’ data originate from distinct generating processes, these methods can be suboptimal and, in some cases, detrimental. Ghosh et al. (2020); Vardhan et al. (2024) address this issue within centralized federated learning by training multiple global models—specifically, $K$ models corresponding to the number of data clusters, with clustering determined by empirical loss to approximate data similarity. Similarly, Sattler et al. (2020) employ cosine similarity on local gradients for clustering clients in FL. Although effective, this leads to a high computational cost as the similarity is computed at stationary solutions of the FL objective. Moreover, there is other notable work on clustered centralized FL which also consider cosine similarity or Euclidean distance between parameters (Briggs et al., 2020; Duan et al., 2021; Ruan & Joe-Wong, 2022).
@@ 1.1 Technical challenges and related work
One notable approach within decentralized learning is gossip learning, a peer-to-peer communication protocol, which has been extensively studied in various machine learning contexts (Kempe et al., 2003; Boyd et al., 2006; Ormándi et al., 2013; Hegedűs et al., 2019). Initial decentralized work using gossip-based optimization for non-convex deep learning tasks, specifically with convolutional neural networks (CNNs), was explored by Blot et al. (2016). However, gossip learning faces challenges in non independent and identically distributed (non-iid) settings where there is a distributional shift across client datasets. This is primarily due to the inadequacy of gossip learning in selectively identifying beneficial peers during the training process. To address these challenges, recent work has addressed decentralized learning within non-iid contexts. Onoszko et al. (2021); Li et al. (2022) introduced client selection algorithms based on empirical loss, employing distinct discovery and selections phases. These studies address challenges in decentralized learning within non-iid settings but require extensive hyperparameter tuning, particularly regarding search length, and rely on hard cluster assignments for clients. Such hard clustering can be detrimental if the discovery phase is noisy. To mitigate these issues, Listo Zec et al. (2022) proposed a sampling strategy that converts the similarity scores into sampling probabilities, which are used each communication round to sample collaborators. This approach identifies similar clients and facilitates model merging in proportion to their similarity. By adopting a soft clustering technique, it enhances the efficiency of peer collaboration, circumventing the need for the computationally intensive and difficult-to-tune discovery phase required by hard clustering methods. Additionally, Listo Zec et al. (2024) introduced a framework using multi-armed bandits for client selection, where the selection process is modeled on empirical rewards derived from validation accuracy.
Quasi-Global Momentum (QGM) (Lin et al., 2021), Neighbourhood Gradient Mean (NGM) (Aketi et al., 2023) and Data-Heterogeneity-Aware Mixing (DHM) (Dandi et al., 2022) are contemporary approaches that address the challenges posed by non-iid data in decentralized learning. While the methods focus on improving consensus optimization, our work diverges significantly in its objectives and assumptions. Specifically, we aim to tackle the problem of varying data distributions across clusters rather than optimizing for a shared consensus model.
@@ 2 Problem formulation
Given that the true distribution $P(x,y)$ is unobserved, direct optimization of equation 1 is not possible. Instead, each client performs empirical risk minimization (ERM) locally by minimizing
$\hat{R}(w_{i})=\mathbb{E}_{(x,y)\sim\mathcal{D}_{i}}\left[\ell(f_{i}(w_{i};x),y)\right],$ (2)
to obtain $w^{\star}_{i}=\arg\min_{w_{i}\in\mathbb{R}^{d}}\hat{R}(w_{i})$.
@@ 2.1 Similarity based peer selection
Algorithm 1 Decentralized learning protocol for a client $i$
1:  Input: model parameters $w_{i}$, temperature $\tau$, learning rate $\eta$, dataset $x_{i},y_{i}\sim\mathcal{D}_{i}$.
2:  Initialize prior probabilities: $p_{ij}=\frac{1}{K-1}$
3:  for each communication round $t=1,\dots,T$ do
4:     Select $m\leq K-1$ clients $\mathcal{M}^{t}_{i}=\{c_{1},c_{2},\dots,c_{m}\}$ without replacement, according to the probability distribution $\mathbf{p}_{i}^{t}=[p^{t}_{i,1},p^{t}_{i,2},\dots,p^{t}_{i,K-1}]$.
5:     Compute similarity scores $s_{ij}^{t}$ between client $i$ and each sampled client $j\in\mathcal{M}^{t}_{i}$.
6:     Update probability vector via softmax: $\mathbf{p}_{i}^{t}\leftarrow\texttt{SOFTMAX}(\mathbf{s}_{i}^{t},\tau)$.
7:     For FedSim, normalize the similarities: $\mathbf{\hat{s}}_{i}=\frac{\mathbf{s_{i}}}{\sum_{j}s_{ij}}$
8:     Aggregate $w_{i}^{t}$ with the models from the sampled clients:
$w_{i}^{\text{merged}}\leftarrow\sum_{j=1}^{m+1}\alpha_{j}w_{j},$
where $\alpha_{j}=\frac{n_{j}}{N}$ for FedAvg, and $\alpha_{j}=\hat{s}_{j}$ for FedSim.
9:     Update the merged model locally for $E$ epochs:
$w_{i}^{t+1}\leftarrow w_{i}^{\text{merged}}-\eta\nabla_{w_{i}}\ell(w_{i}^{\text{merged}};x_{i},y_{i})$
10:  end for
11:  return  Final model parameters for client $i$: $w^{T}_{i}$
@@ 3 Experimental setup
In this section, we describe our experimental setup for studying the effect of common similarity metrics on client identification within decentralized learning. Empirical loss and cosine similarity on gradients have been previously employed (Onoszko et al., 2021; Li et al., 2022; Listo Zec et al., 2022). Additionally, we introduce the Euclidean distance ($L^{2}$) between model parameters and cosine similarity between parameters (as opposed to gradients) as two additional baselines, which have not been tested before in DL for client identification. We define the similarity metrics between two clients $i$ and $j$ as follows:
@@
3.1 Computational cost analysis
When selecting similarity metrics for client identification in decentralized learning, it is crucial to consider their computational overhead, especially in resource-constrained decentralized settings. In this section, we analyze the computational complexity of each of the four similarity metrics at each communication round, focusing on the operations required to compute the similarity score itself, rather than the cost of training epochs.
Inverse empirical loss.
To compute the inverse empirical loss as a similarity metric between client $i$ and client $j$, we evaluate client $i$’s model with parameters $w_{i}$ on every data point $(x,y)$ in client $j$’s local dataset $\mathcal{D}_{j}$. This necessitates performing a forward pass through the model for each data point in $\mathcal{D}_{j}$, resulting in a computational complexity of $\mathcal{O}(N_{j}\times C_{f})$, where $N_{j}$ is the size of the dataset $\mathcal{D}_{j}$ and $C_{f}$ is the cost of a single forward pass. This approach, which deterministically evaluates the loss over the entire dataset $\mathcal{D}_{j}$, imposes a significant computational burden, especially with large datasets or complex models, making it less feasible in practice. While empirical loss could be stochastically estimated using a subset of $\mathcal{D}_{j}$ (e.g., a mini-batch), we use the full dataset to obtain a more deterministic and potentially more reliable measure of functional similarity for guiding peer selection.
Inverse $L^{2}$ distance.
The inverse Euclidean distance between model parameters $w_{i}$ and $w_{j}$ is computed directly from the parameter vectors and has a computational complexity of $\mathcal{O}(P)$, where $P$ is the number of model parameters. This calculation is deterministic and scales linearly with model size, remaining feasible even for large models.
Cosine similarity of weights and gradients.
Both the cosine similarity of weights and the cosine similarity of gradients, as defined by equation 5 using $g^{t}_{i}=w^{t}_{i}-w^{0}_{i}$, have a computational complexity of $\mathcal{O}(P)$, similar to the inverse $L^{2}$ distance. These metrics involve computing dot products and norms of vectors of size $P$, which are deterministic operations manageable even for models with a large number of parameters.
Practical example.
To illustrate the practical implications of these computational costs, consider a moderately sized CNN with $P=1{,}000{,}000$ parameters and a local dataset size of $N_{j}=10{,}000$ samples per client. Assuming an idealized scenario with efficient hardware where the cost of a single forward pass $C_{f}$ is approximately $10$ milliseconds, computing the inverse empirical loss for similarity assessment would require a total time of $N_{j}\times C_{f}=10,000\times 10\;\text{ms}=100,000\;\text{ms}=100\;\text{s}$. This calculation focuses solely on computation time and does not account for potential communication overhead in a decentralized setting.
In contrast, computing the inverse $L^{2}$ distance or cosine similarities involves approximately $P=1{,}000{,}000$ basic vector operations (dot products, norms, subtractions). Assuming each operation takes approximately $1$ nanosecond, the total time is roughly $P\times 1\;\text{ns}=1{,}000{,}000\;\text{ns}=1\;\text{ms}$. This example, while simplified, highlights the order of magnitude difference in computational cost between empirical loss and the other parameter-based similarity metrics, emphasizing the trade-offs between functional similarity measures and computational efficiency.
@@ 3.3 Model selection and hyperparameter tuning
In both federated and decentralized learning environments, the process of model selection and the setting of hyperparameters are crucial yet often under-discussed elements. A common challenge in the literature is the lack of detailed reporting on hyperparameter choices in decentralized settings, which complicates the task of conducting fair and consistent methodological comparisons. Variations in results that stem from differing hyperparameter tuning practices can be mistakenly attributed to the intrinsic qualities of the algorithms themselves, thereby obscuring accurate evaluations. In practice, while some studies use a global validation set, others may employ a local validation set specific to each client, and still others might not clearly disclose their hyperparameter selection strategy.
In our investigation, we strive for a rigorous comparison across all models and similarity metrics by using a local validation set for each client. We tune all hyperparameters using these validation sets, and employ early stopping on each client locally. During communication, some clients may reach early stopping before others. When this occurs, these clients continue to broadcast their models to other clients but cease updating their own models through further training or merging. This approach ensures that our evaluation is both relevant and tailored to the specific conditions of each client’s data. Additionally, we introduce a comparison against an Oracle baseline, which has access to clients with the same data distribution and only communicates with those, providing a benchmark under ideal conditions, and a Local baseline which only performs local training of the available data on a single client.
@@ 3.4 Distributional shifts and datasets
Label shift. Using the CIFAR-10 dataset (Krizhevsky et al., 2009), we create two more realistic and complex scenarios. The first scenario divides the dataset into two clusters based on content categories: an animal cluster, comprising four labels, and a vehicle cluster, with six labels. These clusters consist of 40 and 60 clients, respectively. In the second scenario, we create five equally large clusters of 20 clients where each cluster is defined by two unique, randomly sampled labels. In this way, each cluster has a unique data generating process where $P(y)$ varies while $P(x|y)$ remains constant. We also study large scale pre-trained models under label shift. For this, we use the CIFAR-100 dataset (Krizhevsky et al., 2009), which we split into three distinct clusters of sizes 26, 13 and 13 based on the superclasses.
@@ 3.5 FedSim: leveraging similarity metrics for weighted averaging
Limitations of FedAvg under distribution shift
The FedAvg aggregation does not yield an unbiased estimate of the target distribution $P_{k}$ that a client associated with $P_{k}$ aims to learn when merging models trained on other distribution $P_{i}\neq P_{k}$. Specifically, the expected value of the aggregated model using FedAvg is
$\mathbb{E}[w^{t+1}]=\sum_{i=1}^{m}\frac{N_{i}}{N}\mathbb{E}[w^{t}_{i}]=\sum_{i=1}^{m}\frac{N_{i}}{N}w_{i}^{\star}\neq w_{k}^{\star},$ (8)
where $N_{i}$ is the number of samples for client $i$ and $N=\sum_{i}N_{i}$, $w_{i}^{\star}$ is the optimal model parameters for distribution $P_{i}$, and $w_{k}^{\star}$ the optimal model parameters for distribution $P_{k}$. The last inequality in equation 8 arises because the combination of the optimal $w_{i}^{\star}$ from other distributions $P_{i}$ for $i=1,\dots,m$ does not, in general, equal $w^{\star}_{k}$.
To address this limitation, we propose using similarity-based averaging. Under the assumption that the target distribution $P_{k}$ can be expressed as a weighted combination of the sampled client distributions $P_{i}$, we can show that similarity-based averaging provides and unbiased estimate of the optimal parameters for $P_{k}$.
Assumption 1.
The target $P_{k}(X,Y)$ can be expressed as a weighted combination of the sampled client distributions, i.e. $P_{k}(X,Y)=\sum_{i=1}^{m}s_{i}P_{i}(X,Y)$, where $\mathbf{s}=(s_{1},s_{2},\dots,s_{m})$ is a vector of weights satisfying $s_{i}\geq 0$ and $\sum_{i=1}^{m}s_{i}=1$.
Under Assumption 1, we have the following proposition.
Proposition 1 (Unbiased estimator).
Consider a single round $t$ in the batch stochastic gradient setting with learning rate $\eta$. Let each sampled client $i\in[m]$ compute its local update $w_{i,t}=w_{t}-\eta\nabla_{w}\hat{R}_{i}(w_{t})$, where $\hat{R}_{i}(w_{t})$ is the empirical risk on client $i$. Then, the aggregated update using the similarity weights $\mathbf{s}$ is $w_{t+1}=\sum_{i=1}^{m}s_{i}w_{i,t}$. Under Assumption 1, the expected aggregated update satisfies
$\mathbb{E}[w_{t+1}\mid w_{t}]=\mathbb{E}[w_{t+1}^{P_{k}}\mid w_{t}],$
where $w_{t+1}^{P_{k}}=w_{t}-\eta\nabla_{w}\hat{R}_{P_{k}}(w_{t})$ is the batch SGD update that would be obtained using a sample from the target distribution $P_{k}$.
Proof sketch.
Under Assumption 1, the expected gradient of the aggregated update can be written as
$\mathbb{E}[w_{t+1}\mid w_{t}]$ $=\sum_{i=1}^{m}s_{i}\mathbb{E}[w_{i,t}\mid w_{t}]=\sum_{i=1}^{m}s_{i}\left(w_{t}-\eta\mathbb{E}[\nabla_{w}\hat{R}_{i}(w_{t})]\right)=$ (9)
$w_{t}-\eta\sum_{i=1}^{m}s_{i}\mathbb{E}[\nabla_{w}\hat{R}_{i}(w_{t})]=w_{t}-\eta\mathbb{E}_{(X,Y)\sim P_{k}}[\nabla_{w}\ell(w_{t};X,Y)]=w_{t+1}^{P_{k}}.$
Under this proposition, the expected value of the model update in FedSim with perfect similarity scores $s_{i}$ is
$\mathbb{E}[w^{t+1}]=\sum_{i=1}^{m}s_{i}\mathbb{E}[w_{i}^{t}]=\sum_{i=1}^{m}s_{i}w_{i}^{\star}.$ (10)
Thus, similarity-based averaging with perfect similarity scores provides an unbiased estimate of the optimal model parameters for the target distribution. We present results from our experimental work analyzing common similarity metrics used in decentralized learning in Section 4, both for sampling clients and for aggregation using FedSim.
@@ 4 Results
$L^{2}$ FedAvg $21.10\pm 2.51$ $59.44\pm 0.66$ $84.04\pm 0.63$ $90.49\pm 0.12$ $84.17\pm 0.21$ $\mathbf{55.53\pm 0.57}$
$L^{2}$ FedSim $10.85\pm 0.33$ $57.63\pm 1.14$ $84.06\pm 0.31$ $90.80\pm 0.11$ $84.16\pm 0.01$
Inv loss FedAvg $31.69\pm 3.23$ $\mathbf{61.62\pm 0.83}$ $84.94\pm 0.53$ $89.83\pm 0.22$ $84.83\pm 0.14$ $54.23\pm 0.21$
Inv loss FedSim $14.82\pm 0.33$ $61.08\pm 1.15$ $85.80\pm 0.20$ $90.67\pm 0.07$ $\mathbf{85.08\pm 0.26}$
Cos grad FedAvg $10.32\pm 0.23$ $58.81\pm 0.88$ $\mathbf{86.22\pm 0.17}$ $\mathbf{91.14\pm 0.17}$ $84.87\pm 0.29$ $53.87\pm 0.72$
Cos grad FedSim $10.30\pm 0.25$ $58.13\pm 0.75$ $86.00\pm 0.32$ $90.72\pm 0.24$ $83.32\pm 0.58$
Cos weight FedAvg $10.34\pm 0.28$ $59.14\pm 1.62$ $85.85\pm 0.20$ $91.03\pm 0.23$ $85.04\pm 0.13$ $54.54\pm 0.04$
Cos weight FedSim $\mathbf{10.30\pm 0.21}$ $59.66\pm 1.02$ $85.66\pm 0.36$ $90.62\pm 0.17$ $84.82\pm 0.42$
Random $1494.84\pm 1973$ $59.71\pm 1.13$ $82.73\pm 0.33$ $90.53\pm 0.05$ $83.70\pm 0.12$ $54.88\pm 0.27$
Local $30.26\pm 1.73$ $37.77\pm 0.24$ $77.08\pm 0.98$ $82.22\pm 0.30$ $76.26\pm 0.12$ $29.49\pm 0.03$
@@ 6 Future work
Additionally, investigating the intersection of similarity metrics and privacy is a compelling direction. Beyond model-based similarity metrics like inverse empirical loss, Euclidean distance, and cosine similarity – which assess similarity by comparing clients’ model weights – optimal transport theory offers a framework for measuring similarity based on clients’ data distributions (Peyré et al., 2019). The Wasserstein distance, a fundamental concept in optimal transport, quantifies the minimal cost of transforming one probability distribution into another. While the Wasserstein distance provides a nuanced and theoretically sound measure of data distribution similarity, its practical application is hindered by computational complexity. Computing the exact Wasserstein distance for high-dimensional empirical distributions involves solving optimization problems with significant resource demands. Moreover, using the Wasserstein distance in decentralized learning introduces considerable privacy concerns. Since this metric relies on detailed statistical information or even direct access to empirical data distributions, it risks exposing sensitive information, thereby conflicting with the privacy-preserving goals of decentralized systems. Future work should focus on addressing these computational and privacy challenges associated with advanced similarity metrics like the Wasserstein distance. Developing efficient approximations or privacy-enhancing techniques tailored to these metrics could significantly enhance their utility in decentralized learning, fostering more robust, generalizable, and privacy-conscious machine learning models.